OpenStudy (anonymous):

solve by factoring 6t^2+t-5=0

OpenStudy (blurbendy):

(6t-5)( + ) that will give you a good start

OpenStudy (kirbykirby):

When there's a constant outside the x^2 terms (or t^2) in this case... a good trick is to multiply the constants of the 1st and last term, and think of which two factors, when added, give the constant of the middle term. Here: 6*-5 = -30 conatnat of middle term +t is 1 so -30 = 6*(-5), and 6+(-5) =1 So... we write it it as 6t^2 + 6t - 5t -5 = 0 Then it's clear how to factor this.

OpenStudy (kirbykirby):

Seeing it that way is analogous to factoring when you don't have a coefficient in front of x^2

OpenStudy (anonymous):

\[6 t^2+t-5=(t+1) (6 t-5) \]